OpenStudy (anonymous):

will fan and medal please help me! Tania is 15 years older than Robert. In five years, Tania will be twice as old as Robert is now. How old is Tania right now? a. 30 b. 35 c. 40 d. 45

OpenStudy (mckennarian):

B. 35

OpenStudy (mckennarian):

In five years she will be 40, which will be twice as old as twenty, which is how old robert is.

OpenStudy (anonymous):

Good job I was way too slow to answer that

OpenStudy (mokeira):

Let robert be x in 5 years, Tanya= x+20 x+20=2x x=20 Right now, tanya=x+15 20+15=35

OpenStudy (anonymous):

so the answer is 40?

OpenStudy (mokeira):

35

OpenStudy (anonymous):

oh my bad its B. 35 thanks mckenna

OpenStudy (mckennarian):

No problem